OpenStudy (anonymous):

Solve for x: −3|x − 3| = −6

OpenStudy (anonymous):

x= -5

OpenStudy (anonymous):

Here's a step by step: 1. divide both side by -3 to get \[|x-3| = 2\] 2. break it into to parts: \[x - 3 = 2\] and \[x - 3 = -2\] the reason why is because the absolute value has to possible solutions: positive or negative. 3. solve for x.